# Haitao Xu    Jingru Zhang

### v1

@@ 3.1 Proofs of Lemmas
The first problem aims to decide for any given hinge node or $G$-node $u$, whether a center must be placed at $G_{u}$, and which hanging subtrees of $u$ on $T$ must contain peripheral-center edge or cycles. The following lemma gives the result.
@@
Given a hinge or $G$-node $u$ on $T$, we can decide in $O(mn)$ time whether a center must be placed at point $G_{u}$, and if not, which hanging subtree of $u$ contains a peripheral-center edge or cycle.
@@
For each split subgraph $G_{i}$ of $G_{u}$, let $\mathcal{P}_{i}$ be the subset of all uncertain points whose probability sums in $G_{i}$ are no less than $0.5$ but less than $0.5$ in every other split subgraph. Let $\mathcal{P}^{\prime}$ be the set of all uncertain points whose probability sums in $G_{u}$’s each split subgraph are less than $0.5$, and $\mathcal{P}^{\prime\prime}$ denotes the set of uncertain points that occur in two such subsets $\mathcal{P}^{=}(G_{i})$. Implied by Lemma 2, each $P_{j}\in\mathcal{P}_{i}$ has its median in $H_{i}$ rather than any other hanging subgraph except for $G_{u}$; each $P_{i}\in\mathcal{P}^{\prime}$ has its median exactly at $G_{u}$; each $P_{i}\in\mathcal{P}^{\prime\prime}$ has its median at $G_{u}$ but may also achieve its minimum expected distance at other points in the two split subgraphs where its probability sums are $0.5$.
@@
Lemmas 2 and 3 imply the following properties. If $\Lambda(\mathcal{P}^{\prime},G_{u})>\lambda$ or $\Lambda(\mathcal{P}^{\prime\prime},G_{u})>\lambda$ then $\lambda$ is infeasible due to $\lambda^{*}\geq\max_{P_{k}\in\mathcal{P}}w_{k}\mathsf{E}\mathrm{d}(P_{k},G_{u})$. When $\Lambda(\mathcal{P}^{\prime},G_{u})=\lambda$ or $\Lambda(\mathcal{P}_{j},G_{u})=\lambda$ for some $1\leq j\leq s$, a center must be placed at $G_{u}$ since it covers $\mathcal{P}_{j}$ and most uncertain points not in $\mathcal{P}_{j}$. Moreover, for any $1\leq i\leq s$, if $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, a center must be placed in subgraph $H_{i}/\{h_{i}\}$ so $H_{i}$ contains a peripheral-center edge or cycle; whereas it is not necessary to place a center in $H_{i}/\{h_{i}\}$ if $\Lambda(\mathcal{P}_{i},G_{u})<\lambda$.
@@
Recall that for all $1\leq i\leq s$, subsets $\mathcal{P}^{>}(G_{i})$ and $\mathcal{P}^{=}(G_{i})$ can be computed in $O(mn)$ time. Because subset $\mathcal{P}_{i}$ is exactly the subset $\mathcal{P}^{>}(G_{i})\cup\mathcal{P}^{=}(G_{i})$ excluding all uncertain points that appear in $\mathcal{P}^{=}(G_{i})$ and another subset $\mathcal{P}^{=}(G_{j})$ with $1\leq i\neq j\leq s$. Every uncertain point has its probability sum equal to $0.5$ in at most two split subgraphs of $G_{u}$. Hence, subsets $\mathcal{P}_{i}$ for all $1\leq i\leq s$ can be obtained in $O(mn)$ time.
@@
Additionally, due to $\mathcal{P}^{\prime}=\mathcal{P}-\cup_{i=1}^{s}\{\mathcal{P}^{>}(G_{i})\cup\mathcal{P}^{=}(G_{i})\}$, $\mathcal{P}^{\prime}$ can be known in $O(n)$ time. So is $\mathcal{P}^{\prime\prime}$ on account of $\mathcal{P}^{\prime\prime}=\mathcal{P}-\mathcal{P}^{\prime}-\cup_{i=1}^{s}\mathcal{P}_{i}$. After computing $\mathsf{E}\mathrm{d}(P_{k},G_{u})$ for all $1\leq k\leq n$ in $O(mn)$ time, values $\Lambda(\mathcal{P}^{\prime},G_{u})$, $\Lambda(\mathcal{P}^{\prime\prime},G_{u})$, and $\Lambda(\mathcal{P}_{i},G_{u})$ for each $1\leq i\leq s$ can be obtained in $O(mn)$ time.
@@
Further, based on the above properties, we deal with each of the following cases accordingly. If $\Lambda(\mathcal{P}^{\prime},G_{u})>\lambda$, $\Lambda(\mathcal{P}^{\prime\prime},G_{u})>\lambda$, or more than two split subtrees have $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, then $\lambda$ is not feasible. When $\Lambda(\mathcal{P}^{\prime},G_{u})=\lambda$ or some split subgraph $G_{j}$ has $\Lambda(\mathcal{P}_{j},G_{u})=\lambda$, a center must be placed at $G_{u}$ so we return node $u$. When exactly two split subgraphs have $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, we return the two adjacent nodes of $u$ in its two hanging subtrees respectively representing the two split subgraphs. Similarly, for the case where only one split subgraph has $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, we return $u$’s adjacent node in the corresponding split subtree. Otherwise, $\Lambda(\mathcal{P}^{\prime},G_{u})<\lambda$, $\Lambda(\mathcal{P}^{\prime\prime},G_{u})\leq\lambda$, and $\Lambda(\mathcal{P}_{i},G_{u})<\lambda$ for each $1\leq i\leq s$. Clearly, a center at $G_{u}$ can cover all uncertain points in $\mathcal{P}$ under $\lambda$ so $\lambda$ is feasible.
@@
It is easy to see that the total running time is $O(mn)$. Thus, the lemma is proved. ∎
@@
Similarly, we can compute in $O(mn)$ time the subset $\mathcal{P}_{i}$ for every split subgraph $G_{i}$ of $T^{H}_{u}$. Values $\Lambda(\mathcal{P}_{i},h_{i})$ for all $1\leq i\leq s$ are obtained in the same time complexity since $d(x,h_{i})$ for any $1\leq i\leq s$ is known in $O(1)$ time after an $O(mn)$-time preprocessing work.
By Lemma 3, we have the following observation. For any $1\leq i\leq s$, if $\Lambda(\mathcal{P}_{i},h_{i})>\lambda$ then the hanging subgraph $H_{i}$ must contain a peripheral-center edge or cycle; if $\Lambda(\mathcal{P}_{i},h_{i})=\lambda$ then a center must be placed at hinge $h_{i}$ to cover most uncertain points. Otherwise, $H_{i}$ does not contain any peripheral-center edge or cycle.
@@
For the case where only one split subgraph, e.g., $G_{i}$, has $\Lambda(\mathcal{P}_{i},h_{i})>\lambda$, we apply Lemma 5 to $h_{i}$ in order to decide whether $G_{u}$ contains a peripheral-center edge or cycle. In general, we obtain only node $u$, the node $u^{\prime}$ that is adjacent to the hinge node of $h_{i}$ in $T^{H}_{u}$’s split subtree for $G_{i}$, or both of them. By Lemma 5, the first case means that $G_{u}$ contains all necessary centers so we return only $u$. For the second case, all necessary centers are in $H_{i}$ so $u^{\prime}$ is returned. In the last case, $G_{u}$ is a peripheral-center cycle and $H_{i}$ contains another peripheral-center edge or cycle, so we return both $u$ and $u^{\prime}$.
@@
Let $e$ be the given peripheral-center (out-of-cycle) edge. Assume that it is incident to vertices $v_{1}$ and $v_{2}$ on $T$, which are non-cycle nodes. $e$ has exactly two hanging subgraphs $H_{1},H_{2}$. Suppose $v_{1}$ is in $H_{1}$ and $v_{2}$ is in $H_{2}$. By the definition, a center must be placed on $e$ to cover all uncertain points whose medians are in $H_{1}$ or $H_{2}$.
Recall that subset $\mathcal{P}^{>}(H_{1})\cup\mathcal{P}^{=}(H_{1})$ (resp., $\mathcal{P}^{>}(H_{2})\cup\mathcal{P}^{=}(H_{2})$) is the subset of all uncertain points whose medians are in $H_{1}$ (resp., $H_{2}$), which can be computed in $O(mn)$ time. For convenience, let $\mathcal{P}_{1}=\mathcal{P}^{>}(H_{1})\cup\mathcal{P}^{=}(H_{1})$ and $\mathcal{P}_{2}=\mathcal{P}^{>}(H_{2})\cup\mathcal{P}^{=}(H_{2})$.
To place the necessary center on $e$, we first find the point on $e$ furthest to $v_{1}$ that covers subset $\mathcal{P}_{1}$. If it exists, then we set the weights of all uncertain points covered by it as zero and subsequently solve the one-center problem w.r.t. $\mathcal{P}$ on $G$. Also, we compute such a point on $e$ covering subset $\mathcal{P}_{2}$ and if it exists, solve the one-center problem on $G$ w.r.t. the subset of all uncovered uncertain points. As long as one obtained one-center objective value is no greater than $\lambda$, $\lambda$ is feasible. Otherwise, $\lambda<\lambda^{*}$.
@@
Suppose $v_{1},v_{2},\cdots,v_{s}$ are the clockwise enumeration of all vertices on $G_{u}$. Let $x$ be any point on $G_{u}$. For each $P_{i}\in\mathcal{P}$, consider function $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ in the $x,y$-coordinate plane by setting vertices $v_{1},v_{2},\cdots,v_{s},v_{1}$ on the $x$-axis in order. As analyzed in [9], $\mathsf{E}\mathrm{d}(P_{i},x)$ is a piece-wise linear function of complexity $O(m)$. Hence, solving $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)\leq\lambda$ generates a set $S^{i}$ of $O(m)$ disjoint line segments $s^{i}_{1},s^{i}_{2},\cdots,s^{i}_{z}$ on the $x$-axis, that is, $O(m)$ disjoint arcs on $G_{u}$.
For all $1\leq i\leq n$, functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ in $x\in G_{u}$ can be determined in $O(mn\log mn)$ time [9]. It follows that all $n$ (ordered) sets $S^{i}$ can be built in $O(mn\log mn)$ time.
@@
By the definition, at least one center must be placed on $G_{u_{1}}$ to cover all uncertain points whose medians are in $G^{\prime}$ but that cannot be covered by any points out of $G_{1}$. Recall that subset $\mathcal{P}^{\prime}=\mathcal{P}^{>}(G^{\prime})\cup\mathcal{P}^{=}(G^{\prime})$ is the subset of all uncertain points whose medians are in $G^{\prime}$, i.e., whose probability sums on $G^{\prime}$ are no less than $0.5$. Denote by $\mathcal{P}^{\prime\prime}$ the subset of $\mathcal{P}^{\prime}$ where all uncertain points can be covered by $h_{2}\in G_{u_{2}}$. Clearly, subsets $\mathcal{P}^{\prime}$ and $\mathcal{P}^{\prime\prime}$ can be obtained in $O(mn)$ time.
@@ 4 Computing $\lambda^{*}$
In this section, we present our algorithm that computes $\lambda^{*}$. By the definition, the cycle or out-of-cycle edge on $G$ that contains centers $q^{*}_{1}$ or $q^{*}_{2}$ is a peripheral-center edge or cycle on $G$ under $\lambda^{*}$. We call the cycle or the out-of-cycle edge containing $q^{*}_{1}$ or $q^{*}_{2}$ a critical node or edge. Denote by $c^{*}_{1}$ and $c^{*}_{2}$ the critical edges or cycles.
Similar to the decision algorithm, we perform a ‘binary search’ on $T$ to recursively search for $c^{*}_{1}$ and then for $c^{*}_{2}$ in turn with the support of the following Lemmas 10 and 11.
In short, to find $c^{*}_{1}$, we recursively compute the centroid $c$ of $T$, apply one of Lemmas 10 and 11 to it accordingly, and prune all its split subtrees that are not relevant to $c^{*}_{1}$. After $O(\log mn)$ recursive steps, in general, we obtain an out-of-cycle edge, an edge incident to a cycle node, or a cycle node. For the middle case, that cycle node is exactly the critical node sought. After then, $c^{*}_{2}$ can be found in a similar way by recursively performing the searching on the hanging subtree of the centroid containing $c^{*}_{2}$. Once $c^{*}_{1}$ and $c^{*}_{2}$ are found, Lemma 12 are applied to compute $\lambda^{*}$ with assistance of our decision algorithm.

### v2

@@ 3.1 Proofs of Lemmas
The first problem aims to decide for any given hinge node or $G$-node $u$, whether a center must be placed at vertex $G_{u}$, and which hanging subtrees of $u$ in $T$ must contain peripheral-center edges or cycles. The following lemma gives the result.
@@
Given a hinge or $G$-node $u$ on $T$, we can decide in $O(mn)$ time whether a center must be placed at vertex $G_{u}$, and if not, which hanging subtree of $u$ contains a peripheral-center edge or cycle.
@@
For each split subgraph $G_{i}$ of $G_{u}$, let $\mathcal{P}_{i}$ be the subset of all uncertain points whose probability sums in $G_{i}$ are at least $0.5$ but less than $0.5$ in every other split subgraph. Let $\mathcal{P}^{\prime}$ be the set of all uncertain points whose probability sums in $G_{u}$’s each split subgraph are less than $0.5$, and $\mathcal{P}^{\prime\prime}$ denotes the set of uncertain points that occur in two such subsets $\mathcal{P}^{=}(G_{i})$. Implied by Lemma 2, each $P_{j}\in\mathcal{P}_{i}$ has its median in $H_{i}$ but not in any other hanging subgraph except for $G_{u}$; each $P_{i}\in\mathcal{P}^{\prime}$ has its median (only) at $G_{u}$; each $P_{i}\in\mathcal{P}^{\prime\prime}$ has its median at $G_{u}$ but may also achieve its minimum expected distance at other points in the two split subgraphs where its probability sums are $0.5$.
@@
Lemmas 2 and 3 imply the following properties. If $\Lambda(\mathcal{P}^{\prime},G_{u})>\lambda$ or $\Lambda(\mathcal{P}^{\prime\prime},G_{u})>\lambda$ then $\lambda$ is infeasible due to $\lambda^{*}\geq\max_{P_{k}\in\mathcal{P}}w_{k}\mathsf{E}\mathrm{d}(P_{k},x^{*}_{k})$. When $\Lambda(\mathcal{P}^{\prime},G_{u})=\lambda$ or $\Lambda(\mathcal{P}_{j},G_{u})=\lambda$ for some $1\leq j\leq s$, a center must be placed at $G_{u}$ to cover $\mathcal{P}^{\prime}$ or $\mathcal{P}_{j}$ and most uncertain points in $\mathcal{P}-\mathcal{P}_{j}$. Moreover, for any $1\leq i\leq s$, if $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, center(s) must be placed in subgraph $H_{i}/\{h_{i}\}$ for covering $\mathcal{P}_{i}$, so $H_{i}$ contains peripheral-center edge(s) or cycle(s); otherwise, it is not necessary to place a center in $H_{i}/\{h_{i}\}$ if $\Lambda(\mathcal{P}_{i},G_{u})<\lambda$.
@@
Recall that for all $1\leq i\leq s$, subsets $\mathcal{P}^{>}(G_{i})$ and $\mathcal{P}^{=}(G_{i})$ can be computed in $O(mn)$ time. Because subset $\mathcal{P}_{i}$ is exactly the subset $\mathcal{P}^{>}(G_{i})\cup\mathcal{P}^{=}(G_{i})$ excluding all uncertain points that appear in $\mathcal{P}^{=}(G_{i})$ and other subset $\mathcal{P}^{=}(G_{j})$ with $1\leq i\neq j\leq s$. Every uncertain point has its probability sum equal to $0.5$ in at most two split subgraphs of $G_{u}$. Hence, subsets $\mathcal{P}_{i}$ for all $1\leq i\leq s$ can be obtained in $O(mn)$ time.
@@
Additionally, due to $\mathcal{P}^{\prime}=\mathcal{P}-\cup_{i=1}^{s}\{\mathcal{P}^{>}(G_{i})\cup\mathcal{P}^{=}(G_{i})\}$, $\mathcal{P}^{\prime}$ can be known in $O(n)$ time. So is $\mathcal{P}^{\prime\prime}$ on account of $\mathcal{P}^{\prime\prime}=\mathcal{P}-\mathcal{P}^{\prime}-\cup_{i=1}^{s}\mathcal{P}_{i}$. Because computing $\mathsf{E}\mathrm{d}(P_{k},G_{u})$ for all $1\leq k\leq n$ takes in $O(mn)$ time [9]. It then follows that values $\Lambda(\mathcal{P}^{\prime},G_{u})$, $\Lambda(\mathcal{P}^{\prime\prime},G_{u})$, and $\Lambda(\mathcal{P}_{i},G_{u})$ for each $1\leq i\leq s$ can be obtained in $O(mn)$ time.
@@
Further, based on the above properties, we deal with each of the following cases accordingly. If $\Lambda(\mathcal{P}^{\prime},G_{u})>\lambda$, $\Lambda(\mathcal{P}^{\prime\prime},G_{u})>\lambda$, or more than two split subtrees have $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, then $\lambda$ is not feasible. When $\Lambda(\mathcal{P}^{\prime},G_{u})=\lambda$ or some split subgraph $G_{j}$ has $\Lambda(\mathcal{P}_{j},G_{u})=\lambda$, a center must be placed at $G_{u}$ so we return node $u$. When exactly two split subgraphs have $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, we return the two adjacent nodes of $u$ in its two hanging subtrees respectively representing the two split subgraphs. Similarly, for the case where only one split subgraph has $\Lambda(\mathcal{P}_{i},G_{u})>\lambda$, we return $u$’s adjacent node in the corresponding split subtree. Otherwise, $\Lambda(\mathcal{P}^{\prime},G_{u})<\lambda$, $\Lambda(\mathcal{P}^{\prime\prime},G_{u})\leq\lambda$, and $\Lambda(\mathcal{P}_{i},G_{u})<\lambda$ for each $1\leq i\leq s$. Clearly, a center at $G_{u}$ can cover all uncertain points in $\mathcal{P}$ under $\lambda$, so $\lambda$ is feasible.
@@
It is clear to see that the total running time is $O(mn)$. Thus, the lemma is proved. ∎
@@
Similarly, we can compute the subset $\mathcal{P}_{i}$ for every split subgraph $G_{i}$ of $T^{H}_{u}$ in $O(mn)$ time. Values $\Lambda(\mathcal{P}_{i},h_{i})$ for all $1\leq i\leq s$ are obtained in the same time complexity since for any point $x$, $d(x,h_{i})$ with $1\leq i\leq s$ can be known in $O(1)$ time after an $O(mn)$-time preprocessing work.
By Lemma 2 and Lemma 3, we have the following observation. For any $1\leq i\leq s$, if $\Lambda(\mathcal{P}_{i},h_{i})>\lambda$ then the hanging subgraph $H_{i}$ must contain a peripheral-center edge or cycle; if $\Lambda(\mathcal{P}_{i},h_{i})=\lambda$ then a center must be placed at hinge $h_{i}$ to cover most uncertain points. Otherwise, $H_{i}$ does not contain any peripheral-center edge or cycle.
@@
For the case where only one split subgraph, e.g., $G_{i}$, has $\Lambda(\mathcal{P}_{i},h_{i})>\lambda$, we apply Lemma 5 to $h_{i}$ in order to decide whether $G_{u}$ is a peripheral-center cycle. In general, we obtain only node $u$, the node $u^{\prime}$ that is adjacent to the hinge node of $h_{i}$ in $T^{H}_{u}$’s split subtree for $G_{i}$, or both of them. By Lemma 5, the first case means that $G_{u}$ contains all necessary centers so we return only $u$. For the second case, all necessary centers are in $H_{i}$ so $u^{\prime}$ is returned. In the last case, $G_{u}$ is a peripheral-center cycle and $H_{i}$ contains another peripheral-center edge or cycle, so we return both $u$ and $u^{\prime}$.
@@
Let $e$ be the given peripheral-center (out-of-cycle) edge. Assume that it is incident to vertices $v_{1}$ and $v_{2}$ on $T$, which are non-cycle nodes. $e$ has exactly two hanging subgraphs $H_{1},H_{2}$. Suppose $v_{1}$ is in $H_{1}$ and $v_{2}$ is in $H_{2}$. By the definition, a center must be placed on $e$ to cover all uncertain points that have medians all in $H_{1}$ or $H_{2}$.
Recall that subset $\mathcal{P}^{>}(H_{1})\cup\mathcal{P}^{=}(H_{1})$ (resp., $\mathcal{P}^{>}(H_{2})\cup\mathcal{P}^{=}(H_{2})$) is the subset of all uncertain points whose medians are in $H_{1}$ (resp., $H_{2}$), which can be obtained in $O(mn)$ time. For convenience, let $\mathcal{P}_{1}=\mathcal{P}^{>}(H_{1})\cup\mathcal{P}^{=}(H_{1})$ and $\mathcal{P}_{2}=\mathcal{P}^{>}(H_{2})\cup\mathcal{P}^{=}(H_{2})$.
To place necessary center(s) on $e$, we first find the point on $e$ furthest to $v_{1}$ that covers subset $\mathcal{P}_{1}$. If it exists, then we set the weights of all uncertain points covered by it as zero and subsequently solve the one-center problem w.r.t. $\mathcal{P}$ on $G$. Also, we compute such a point on $e$ covering subset $\mathcal{P}_{2}$ and if it exists, solve the one-center problem on $G$ w.r.t. the subset of all uncovered uncertain points. As long as one obtained one-center objective value is no greater than $\lambda$, $\lambda$ is feasible. Otherwise, $\lambda<\lambda^{*}$.
@@
Suppose $v_{1},v_{2},\cdots,v_{s}$ are the clockwise enumeration of all vertices on $G_{u}$. Let $x$ be any point on $G_{u}$. For each $P_{i}\in\mathcal{P}$, consider function $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ in the $x,y$-coordinate plane by setting vertices $v_{1},v_{2},\cdots,v_{s},v_{1}$ on the $x$-axis in order. As analyzed in [9], $\mathsf{E}\mathrm{d}(P_{i},x)$ is a piecewise linear function of complexity $O(m)$. Hence, solving $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)\leq\lambda$ generates a set $S^{i}$ of $O(m)$ disjoint line segments $s^{i}_{1},s^{i}_{2},\cdots,s^{i}_{z}$ on the $x$-axis, that is, $O(m)$ disjoint arcs on $G_{u}$.
For all $1\leq i\leq n$, functions $y=w_{i}\mathsf{E}\mathrm{d}(P_{i},x)$ in $x\in G_{u}$ can be determined in $O(mn\log mn)$ time [9]. Because each $y=\mathsf{E}\mathrm{d}(P_{i},x)$ is piecewise linear and of complexity $O(m)$. It follows that all $n$ (ordered) sets $S^{i}$, generated by solving $w_{i}\mathsf{E}\mathrm{d}(P_{i},x)\leq\lambda$, can be built in $O(mn\log mn)$ time.
@@
By the definition, at least one center must be placed on $G_{u_{1}}$ to cover all uncertain points whose medians are in $G^{\prime}$ but that cannot be covered by any points out of $G_{u_{1}}$. Recall that subset $\mathcal{P}^{\prime}=\mathcal{P}^{>}(G^{\prime})\cup\mathcal{P}^{=}(G^{\prime})$ is the subset of all uncertain points whose medians are in $G^{\prime}$, i.e., whose probability sums on $G^{\prime}$ are no less than $0.5$. Denote by $\mathcal{P}^{\prime\prime}$ the subset of $\mathcal{P}^{\prime}$ where all uncertain points can be covered by $h_{2}\in G_{u_{2}}$. Clearly, subsets $\mathcal{P}^{\prime}$ and $\mathcal{P}^{\prime\prime}$ can be obtained in $O(mn)$ time.
@@ 4 Computing $\lambda^{*}$
In this section, we present our algorithm that computes $\lambda^{*}$. By the definitions, the cycle or out-of-cycle edge on $G$ that contains centers $q^{*}_{1}$ or $q^{*}_{2}$ is a peripheral-center edge or cycle on $G$ under $\lambda^{*}$. We call the cycle or the out-of-cycle edge containing $q^{*}_{1}$ or $q^{*}_{2}$ a critical node or edge. Denote by $c^{*}_{1}$ and $c^{*}_{2}$ the critical edges or cycles.
Similar to the decision algorithm, we perform two “binary searches” on $T$ to respectively find $c^{*}_{1}$ and $c^{*}_{2}$ with the support of the following Lemmas 10 and 11.
In short, to find $c^{*}_{1}$, we recursively compute the centroid $c$ of $T$, apply Lemmas 10 or 11 to it accordingly, and prune all its split subtrees that are not relevant to $c^{*}_{1}$. After $O(\log mn)$ recursive steps, in general, we obtain an out-of-cycle edge, an edge incident to a cycle node, or a cycle node. For the middle case, that cycle node is exactly the critical node sought. After then, $c^{*}_{2}$ can be found in a similar way by recursively performing the searching on the hanging subtree of the centroid containing $c^{*}_{2}$. Once $c^{*}_{1}$ and $c^{*}_{2}$ are found, Lemma 12 are applied to compute $\lambda^{*}$ with assistance of our decision algorithm.